\begin{thebibliography}{99}

\bibitem{Whee} R. Ruffini and J. A. Wheeler, Phys. Today {\bf 24}, 30
(1971).

\bibitem{Car} B. Carter, in {\it Black Holes}, Proceedings of 1972 Session of Ecole d'ete de Physique Theorique,
edited by C. De Witt and B. S. De Witt (Gordon and Breach, New York,
1973).

\bibitem{Bekto} J. D. Bekenstein, Physics Today {\bf 33}, 24 (1980).

\bibitem{Cha} J. E. Chase, Commun. Math. Phys. {\bf 19}, 276 (1970).

\bibitem{Bek1} J. D. Bekenstein, Phys. Rev. Letters {\bf 28}, 452 (1972).

\bibitem{Teit} C. Teitelboim, Lett. Nuov. Cim. {\bf 3}, 326 (1972).

\bibitem{Noteit} It is worth mentioning that recent studies \cite{Hodrc,Herkr}
of the composed Einstein-scalar system have explicitly proved that
{\it spinning} black holes can support {\it stationary} (rather than
static) matter configurations made of massive scalar (bosonic)
fields.

\bibitem{Hodrc} S. Hod, Phys. Rev. D {\bf 86}, 104026 (2012) [arXiv:1211.3202];
S. Hod, The Euro. Phys. Journal C {\bf 73}, 2378 (2013)
[arXiv:1311.5298]; S. Hod, Phys. Rev. D {\bf 90}, 024051 (2014)
[arXiv:1406.1179]; S. Hod, Phys. Lett. B {\bf 739}, 196 (2014)
[arXiv:1411.2609]; S. Hod, Class. and Quant. Grav. {\bf 32}, 134002
(2015) [arXiv:1607.00003]; S. Hod, Phys. Lett. B {\bf 751}, 177
(2015); S. Hod, Class. and Quant. Grav. {\bf 33}, 114001 (2016); S.
Hod, Phys. Lett. B {\bf 758}, 181 (2016) [arXiv:1606.02306]; S. Hod
and O. Hod, Phys. Rev. D {\bf 81}, 061502 Rapid communication (2010)
[arXiv:0910.0734]; S. Hod, Phys. Lett. B {\bf 708}, 320 (2012)
[arXiv:1205.1872]; S. Hod, Jour. of High Energy Phys. {\bf 01}, 030
(2017) [arXiv:1612.00014].

\bibitem{Herkr} C. A. R. Herdeiro and E. Radu, Phys. Rev. Lett. {\bf 112}, 221101
(2014); C. L. Benone, L. C. B. Crispino, C. Herdeiro, and E. Radu,
Phys. Rev. D {\bf 90}, 104024 (2014); C. A. R. Herdeiro and E. Radu,
Phys. Rev. D {\bf 89}, 124018 (2014); C. A. R. Herdeiro and E. Radu,
Int. J. Mod. Phys. D {\bf 23}, 1442014 (2014); Y. Brihaye, C.
Herdeiro, and E. Radu, Phys. Lett. B {\bf 739}, 1 (2014); J. C.
Degollado and C. A. R. Herdeiro, Phys. Rev. D {\bf 90}, 065019
(2014); C. Herdeiro, E. Radu, and H. R\'unarsson, Phys. Lett. B {\bf
739}, 302 (2014); C. Herdeiro and E. Radu, Class. Quantum Grav. {\bf
32} 144001 (2015); C. A. R. Herdeiro and E. Radu, Int. J. Mod. Phys.
D {\bf 24}, 1542014 (2015); C. A. R. Herdeiro and E. Radu, Int. J.
Mod. Phys. D {\bf 24}, 1544022 (2015); P. V. P. Cunha, C. A. R.
Herdeiro, E. Radu, and H. F. R\'unarsson, Phys. Rev. Lett. {\bf
115}, 211102 (2015); B. Kleihaus, J. Kunz, and S. Yazadjiev, Phys.
Lett. B {\bf 744}, 406 (2015); C. A. R. Herdeiro, E. Radu, and H. F.
R\'unarsson, Phys. Rev. D {\bf 92}, 084059 (2015); C. Herdeiro, J.
Kunz, E. Radu, and B. Subagyo, Phys. Lett. B {\bf 748}, 30 (2015);
C. A. R. Herdeiro, E. Radu, and H. F. R\'unarsson, Class. Quant.
Grav. {\bf 33}, 154001 (2016); C. A. R. Herdeiro, E. Radu, and H. F.
R\'unarsson, Int. J. Mod. Phys. D {\bf 25}, 1641014 (2016); Y.
Brihaye, C. Herdeiro, and E. Radu, Phys. Lett. B {\bf 760}, 279
(2016); Y. Ni, M. Zhou, A. C. Avendano, C. Bambi, C. A. R. Herdeiro,
and E. Radu, JCAP {\bf 1607}, 049 (2016); M. Wang, arXiv:1606.00811
.

\bibitem{Heu} M. Heusler, J. Math. Phys. {\bf 33}, 3497 (1992);
M. Heusler, Class. Quant. Grav. {\bf 12}, 779 (1995).

\bibitem{Bek2} J. D. Bekenstein, Phys. Rev. D {\bf 51}, R6608 (1995).

\bibitem{BekMay} A. E. Mayo and J. D. Bekenstein, Phys. Rev. D {\bf 54}, 5059 (1996).

\bibitem{Saa} A. Saa, Phys. Rev.D {\bf 53}, 7377 (1996).

\bibitem{Notesa} It is worth mentioning that, for spherically symmetric
$(3+1)$-dimensional black holes, Saa \cite{Saa} has also
demonstrated the absence of scalar hair $\psi(r)$ in the nonminimal
coupling regime $0<\xi<1/6$ with $8\pi\psi^2(r)<1/\xi$ or with
$1/\xi<8\pi\psi^2(r)<[6\xi(1/6-\xi)]^{-1}$ and in the nonminimal
coupling regime $\xi>1/6$ with $8\pi\psi^2(r)\neq 1/\xi$. In the
present paper we shall provide a unified no-hair theorem for
nonminimally coupled scalar fields which is valid for {\it generic}
values of the dimensionless coupling parameter $\xi$ and for {\it
generic} values of the scalar field eigenfunction $\psi(r)$.

\bibitem{Notell} As shown in \cite{BekMay}, the expansion coefficient
is given by $L\equiv 1+8\pi T^{t}_{t}(r_{\text{H}})
r^2_{\text{H}}>0$, where $-T^{t}_{t}(r_{\text{H}})$ is the energy
density of the matter fields at the black-hole horizon.

\bibitem{Abram} M. Abramowitz and I. A. Stegun, {\it Handbook of
Mathematical Functions} (Dover Publications, New York, 1970).

\bibitem{Noteb0} It is worth noting that, for $\beta=0$, one
finds from (\ref{Eq13}) the asymptotic near-horizon functional
behavior $\psi(x\to0)=A\cdot\ln(x)+B$, where $\{A,B\}$ are
constants. Substituting this expression into (\ref{Eq8}) and taking
cognizance of the energy condition (\ref{Eq10}) \cite{BekMay}, one
immediately realizes that, for $A\neq0$, this near-horizon scalar
function is physically unacceptable. One therefore concludes that,
for physically acceptable spacetimes, the scalar eigenfunction is
strictly constant. Substituting $\psi=B$ into Eqs. (\ref{Eq8}) and
(\ref{Eq9}), one finds that the components of the energy momentum
tensor are identically {\it zero}. Thus, the constant $B$ has no
influence on physical quantities and one may therefore take $B=0$
without loss of generality.
%For the special case of $R=0$ ($\beta=0$), one can write the scalar
%radial equation (\ref{Eq7}) in the form
%$[\ln(\psi{'})]{'}=-[2/r+(\nu-\lambda){'}/2]$, which yields the
%exact solution $\psi{'}=A\cdot r^{-2}e^{(\lambda-\nu)/2}$, where $A$
%is a constant. This solution, together with Eq. (\ref{Eq2}) and the
%characteristic near-horizon relation $e^{\nu}=N\cdot x+O(x^2)$ with
%$N>0$ \cite{BekMay} [see also Eq. (\ref{Eq3})], yield the asymptotic
%near-horizon functional behavior $d\psi/dx=A[(LN)^{-1/2}\cdot
%x^{-1}+O(1)]$, which immediately implies
%$\psi(x)=A[(LN)^{-1/2}\cdot\ln(x)+O(x)]+B$, where $B$ is a constant.
%Substituting this expression into (\ref{Eq8}) and taking cognizance
%of the near-horizon behavior (\ref{Eq2}), one realizes that
%physically acceptable scalar solutions that respect the energy
%condition (\ref{Eq10}) \cite{BekMay} must be characterized by $A=0$.
%One therefore concludes that the scalar eigenfunction is strictly
%constant in the entire spacetime. Substituting $\psi=B$ into Eqs.
%(\ref{Eq8}) and (\ref{Eq9}), one finds that the components of the
%energy momentum tensor are identically {\it zero}. Thus, the
%constant $B$ has no influence on physical quantities and one may
%take $B=0$ without loss of generality. We therefore conclude that,
%for the special $R=0$ case, there is no nonminimally coupled
%external scalar hair.

\end{thebibliography}

\documentclass[article,preprint,groupedaddress]{revtex4}
\usepackage{epsfig}
\usepackage{graphicx}
\usepackage{amssymb}

\begin{document}
\title{No nonminimally coupled massless scalar hair for spherically symmetric neutral black holes}
\author{Shahar Hod}
\affiliation{The Ruppin Academic Center, Emeq Hefer 40250, Israel}
\affiliation{ }
\affiliation{The Hadassah Institute, Jerusalem 91010, Israel}
\date{\today}

\begin{abstract}
\ \ \ We provide a remarkably compact proof that spherically
symmetric neutral black holes cannot support static nonminimally
coupled massless scalar fields. The theorem is based on causality
restrictions imposed on the energy-momentum tensor of the fields
near the regular black-hole horizon.

\end{abstract}
\bigskip
\maketitle

\section{Introduction}

The non-linearly coupled Einstein-scalar field equations have
attracted the attention of physicists and mathematicians for more
than five decades. Interestingly, the composed Einstein-scalar
system is characterized by very powerful and elegant no-hair
theorems \cite{Whee,Car,Bekto} which rule out the existence of
asymptotically flat black-hole solutions with regular event horizons
that support various types of scalar (spin-0) matter configurations.

The early no-hair theorems of Chase \cite{Cha} and Bekenstein
\cite{Bek1} have ruled out the existence of regular black holes
supporting static minimally coupled massless scalar field
configurations. The no-hair theorems of Bekenstein \cite{Bek1} and
Teitelboim \cite{Teit} have excluded the existence of black-hole
hair made of minimally coupled massive scalar fields
\cite{Noteit,Hodrc,Herkr}. Later no-hair theorems of Heusler
\cite{Heu} and Bekenstein \cite{Bek2} have ruled out the existence
of neutral black-hole spacetimes supporting static matter
configurations made of minimally coupled scalar fields with positive
semidefinite self-interaction potentials.

The physically interesting regime of scalar fields nonminimally
coupled to gravity has been investigated by several authors. In a
very interesting paper, Mayo and Bekenstein \cite{BekMay} have
proved that spherically symmetric charged black holes cannot support
matter configurations made of charged scalar fields nonminimally
coupled to gravity with generic values of the dimensionless coupling
parameter $\xi$ [The physical parameter $\xi$ quantifies the
nonminimal coupling of the field to gravity, see Eq. (\ref{Eq4})
below]. Intriguingly, the rigorous derivation of a no-hair theorem
for {\it neutral} scalar fields nonminimally coupled to gravity
seems to be a mathematically more challenging task. In particular,
the important no-hair theorems of \cite{BekMay,Saa} can be used to
rule out the existence of spherically symmetric scalar hairy
configurations in the restricted physical regimes $\xi<0$ and
$\xi\geq1/2$ \cite{Notesa}.

The main goal of the present paper is to present a (remarkably
compact) unified no-hair theorem for neutral massless scalar fields
nonminimally coupled to gravity which is valid for {\it generic}
values of the dimensionless coupling parameter $\xi$ (In particular,
below we shall extend the interesting no-scalar hair theorems of
\cite{BekMay} and \cite{Saa} to the physical regime of nonminimally
coupled neutral scalar fields with $0<\xi<1/2$). Our theorem, to be
proved below, is based on simple physical restrictions imposed by
causality on the energy-momentum tensor of the fields near the
regular horizon of the black-hole spacetime.

\section{Description of the system}

We consider a non-linear physical system composed of a neutral black
hole of horizon radius $r_{\text{H}}$ and a massless scalar field
$\psi$ with nonminimal coupling to gravity. The composed
black-hole-scalar-field system is assumed to be static and
spherically symmetric, in which case the spacetime outside the
black-hole horizon is characterized by the curved line element
\cite{BekMay} (We shall use natural units in which $G=c=1$)
\begin{equation}\label{Eq1}
ds^2=-e^{\nu}dt^2+e^{\lambda}dr^2+r^2(d\theta^2+\sin^2\theta
d\phi^2)\ ,
\end{equation}
where $\nu=\nu(r)$ and $\lambda=\lambda(r)$ [Here
$(t,r,\theta,\phi)$ are the Schwarzschild coordinates]. As
explicitly proved in \cite{BekMay}, regardless of the matter content
of the curved spacetime, a non-extremal regular black hole is
characterized by the near-horizon relations \cite{Notell}
\begin{equation}\label{Eq2}
e^{-\lambda}=L\cdot x+O(x^2)\ \ \ \ \text{where}\ \ \ \ x\equiv
{{r-r_{\text{H}}}\over{r_{\text{H}}}}\ \ \ ; \ \ \ L>0\
\end{equation}
and
\begin{equation}\label{Eq3}
\lambda{'}r_{\text{H}}=-{{1}\over{x}}+O(1)\ \ \ \ ; \ \ \ \
\nu{'}r_{\text{H}}={{1}\over{x}}+O(1)\  .
\end{equation}

The curved black-hole spacetime is non-linearly and non-minimally
coupled to a real massless scalar field $\psi$ whose action is given
by \cite{BekMay}
\begin{equation}\label{Eq4}
S=S_{EH}-{1\over2}\int\big(\partial_{\alpha}\psi\partial^{\alpha}\psi+\xi
R\psi^2\big)\sqrt{-g}d^4x\ ,
\end{equation}
where the dimensionless physical parameter $\xi$ quantifies the
strength of the nonminimal coupling of the field to gravity, $R(r)$
is the scalar curvature of the spacetime, and $S_{\text{EH}}$ is the
Einstein-Hilbert action. As explicitly shown in \cite{BekMay}, in
the near-horizon $x\ll1$ region, $R$ is given by the simple
expression
\begin{equation}\label{Eq5}
R={{4L-2}\over{r^2_{\text{H}}}}\cdot[1+O(x)]\  .
\end{equation}

From the action (\ref{Eq4}) one finds the characteristic
differential equation \cite{BekMay}
\begin{equation}\label{Eq6}
\partial_{\alpha}\partial^{\alpha}\psi-\xi R\psi=0\
\end{equation}
for the nonminimally coupled scalar field. Using the metric
components (\ref{Eq1}) of the curved black-hole spacetime, one can
express the scalar radial equation in the form
\begin{equation}\label{Eq7}
\psi{''}+{1\over2}\big({{4}\over{r}}+\nu{'}-\lambda{'}\big)\psi{'}-\xi
R e^{\lambda}\psi=0\  .
\end{equation}
(Here a prime ${'}$ denotes a spatial derivative with respect to the
radial coordinate $r$).

The action (\ref{Eq4}) also yields the compact expressions
\cite{BekMay}
\begin{equation}\label{Eq8}
T^{t}_{t}=e^{-\lambda}{{\xi(4/r-\lambda{'})\psi\psi{'}+(2\xi-1/2)(\psi{'})^2+2\xi\psi\psi{''}}
\over{1-8\pi\xi\psi^2}}\
\end{equation}
and
\begin{equation}\label{Eq9}
T^{t}_{t}-T^{\phi}_{\phi}=e^{-\lambda}{{\xi(2/r-\nu{'})\psi\psi{'}}
\over{1-8\pi\xi\psi^2}}
\end{equation}
for the components of the energy-momentum tensor. As explicitly
proved in \cite{BekMay}, regardless of the matter content of the
theory, a regular hairy black-hole spacetime must be characterized
by finite mixed components of the energy-momentum tensor:
\begin{equation}\label{Eq10}
\{|T^{t}_{t}|,|T^{r}_{r}|,|T^{\theta}_{\theta}|,|T^{\phi}_{\phi}|\}<\infty\
.
\end{equation}
In addition, it was proved in \cite{BekMay} that causality
requirements enforce the characteristic inequalities
\footnotemark[1]
\begin{equation}\label{Eq11}
|T^{\theta}_{\theta}|=|T^{\phi}_{\phi}|\leq|T^{t}_{t}|\geq|T^{r}_{r}|\
\end{equation}
on the components of the energy-momentum tensors of physically
acceptable systems. Note that the relations \cite{BekMay}
\begin{equation}\label{Eq12}
\text{sgn}(T^{t}_{t})=\text{sgn}(T^{t}_{t}-T^{r}_{r})=\text{sgn}(T^{t}_{t}-T^{\phi}_{\phi})\
\end{equation}
provide necessary conditions for the validity of the characteristic
energy conditions (\ref{Eq11}).

\footnotetext[1]{As explicitly shown by Bekenstein and Mayo
\cite{BekMay}, for spherically symmetric spacetimes one can write
$\epsilon=-T^{t}_{t}-\sum_{i=1}^{3}c^2_i(T^{t}_{t}-T^{i}_{i})$ and
$j^{\mu}j_{\mu}=-(T^{t}_{t})^2-\sum_{i=1}^{3}c^2_i[(T^{t}_{t})^2-(T^{i}_{i})^2]$,
where $\epsilon\equiv T_{\mu\nu}u^{\mu}u^{\nu}$ and
$j^{\mu}\equiv-T^{\mu}_{\nu}u^{\nu}$ are respectively the energy
density and the Poynting vector according to a physical observer
with a 4-velocity $u^{\nu}$, and the coefficients
$\{c_i\}_{i=0}^{3}$ are characterized by the normalization condition
$-c^2_0+\sum_{i=1}^{3}c^2_i=-1$ (this relation guarantees that
$u^{\mu}u_{\mu}=-1$ \cite{BekMay}). For physically acceptable
systems in which the transfer of energy is not superluminal, the
energy density should be of the same sign as $-T^{t}_{t}$ and the
Poynting vector should be non-spacelike ($j^{\mu}j_{\mu}\leq0$) for
all observers \cite{BekMay} (that is, for all choices of the
coefficients $\{c_i\}_{i=0}^{3}$). These physical requirements yield
the characteristic energy conditions (\ref{Eq11}) \cite{BekMay}.}

\section{The no-hair theorem for static nonminimally coupled massless scalar fields}

In the present section we shall explicitly prove that a spherically
symmetric non-extremal neutral black hole {\it cannot} support
non-linear hair made of static nonminimally coupled massless scalar
fields.

We start our proof with the scalar radial equation (\ref{Eq7})
which, in the near-horizon $x\ll1$ region, can be written in the
form [see Eqs. (\ref{Eq2}), (\ref{Eq3}), and (\ref{Eq5})]
\begin{equation}\label{Eq13}
{{d^2\psi}\over{dx^2}}+{{1}\over{x}}{{d\psi}\over{dx}}+{{\beta}\over{x}}\psi=0\
\ \ \ ; \ \ \ \ \beta\equiv \xi(2-4L)/L\  .
%\psi{''}+{{1}\over{xr_{\text{H}}}}\psi{'}-{{\beta}\over{x}}\psi=0\ \
%\ \ ; \ \ \ \ \beta\equiv \xi R(r_{\text{H}})/L\  .
\end{equation}
The general mathematical solution of the ordinary differential
equation (\ref{Eq13}) can be expressed in terms of the familiar
Bessel functions (see Eq. 9.1.53 of \cite{Abram})
\begin{equation}\label{Eq14}
\psi(x)=A\cdot J_0(2\beta^{1/2}x^{1/2})+ B\cdot
Y_0(2\beta^{1/2}x^{1/2})\ \ \ \ \text{for}\ \ \ \ x\ll1\  ,
\end{equation}
where $\{A,B\}$ are constants. Using Eqs. 9.1.8 and 9.1.12 of
\cite{Abram}, one finds the asymptotic near-horizon behavior
\begin{equation}\label{Eq15}
\psi(x\to0)=A\cdot[1-\beta x+O(x^2)]+ B\cdot [\pi^{-1}\ln(\beta
x)+O(1)]\
\end{equation}
of the radial scalar function. Substituting Eqs. (\ref{Eq2}),
(\ref{Eq3}), and (\ref{Eq15}) into Eq. (\ref{Eq8}) and taking
cognizance of the energy condition (\ref{Eq10}) \cite{BekMay}, one
immediately realizes that the coefficient of the singular term in
the asymptotic near-horizon solution (\ref{Eq15}) should vanish
\cite{Noteb0}:
\begin{equation}\label{Eq16}
B=0\  .
\end{equation}

We therefore find that the nonminimally coupled scalar field is
characterized by the near-horizon behavior
\begin{equation}\label{Eq17}
\psi(x\ll1)=A\cdot J_0(2\beta^{1/2}x^{1/2})\  .
\end{equation}
Substituting (\ref{Eq17}) into (\ref{Eq8}) and (\ref{Eq9}) and using
the near-horizon relations (\ref{Eq2}) and (\ref{Eq3}), one obtains
the simple expressions
\begin{equation}\label{Eq18}
T^{t}_{t}=\xi\cdot{{L\psi\psi{'}}\over{r_{\text{H}}(1-8\pi\xi\psi^2)}}\cdot[1+O(x)]\
%\ \ \ \text{for}\ \ \ \ x\ll1\
\end{equation}
and
\begin{equation}\label{Eq19}
T^{t}_{t}-T^{\phi}_{\phi}=-\xi\cdot{{L\psi\psi{'}}\over{r_{\text{H}}(1-8\pi\xi\psi^2)}}\cdot[1+O(x)]\
%\ \ \ \text{for}\ \ \ \ x\ll1\
\end{equation}
for the components of the energy-momentum tensor in the near-horizon
$x\ll1$ region. We immediately find from (\ref{Eq18}) and
(\ref{Eq19}) the near-horizon relation
\begin{equation}\label{Eq20}
T^{t}_{t}=-(T^{t}_{t}-T^{\phi}_{\phi})\  ,
\end{equation}
in {\it contradiction} with the characteristic relation (\ref{Eq12})
imposed by causality on the energy-momentum components of physically
acceptable systems.

\section{Summary}

In this compact analysis, we have proved that {\it if} a spherically
symmetric neutral black hole can support non-linear configurations
made of nonminimally coupled massless scalar fields, then in the
near-horizon $(r-r_{\text{H}})/r_{\text{H}}\ll1$ region the energy
momentum components of the fields must be characterized by the
relation $T^{t}_{t}=-(T^{t}_{t}-T^{\phi}_{\phi})$ [see Eq.
(\ref{Eq20})]. However, one realizes that this near-horizon behavior
is in {\it contradiction} with the characteristic relation
$\text{sgn}(T^{t}_{t})=\text{sgn}(T^{t}_{t}-T^{\phi}_{\phi})$ [see
Eq. (\ref{Eq12})] which, as explicitly proved in \cite{BekMay}, is
imposed by causality on the energy-momentum components of generic
physically acceptable systems. Thus, there are no physically
acceptable solutions for the eigenfunction of the external
nonminimally coupled massless scalar fields except the trivial one,
$\psi\equiv0$.

We therefore conclude that spherically symmetric neutral black holes
cannot support static configurations made of nonminimally coupled
massless scalar fields with {\it generic} values of the
dimensionless physical parameter $\xi$.

%\newpage

\bigskip
\noindent
{\bf ACKNOWLEDGMENTS}
\bigskip

This research is supported by the Carmel Science Foundation. I would
like to thank Yael Oren, Arbel M. Ongo, Ayelet B. Lata, and Alona B.
Tea for helpful discussions.

%\newpage

\end{document}